Riva takes 5 hours to do a job. If Megan takes 8 hours and Sarah takes 10 hours to do the same job, then how many hours will it take all of them to do the job if they work together?
Answer by J2R2R(94)

Time
Riva = 5 hours
Megan = 8 hours
Sarah = 10 hours
Speed
Riva = 1/5 of the job per hour
Megan = 1/8 of the job per hour
Sarah = 1/10 of the job per hour

Combine all speeds
All = 1/5 + 1/8 + 1/10 = 8/40 + 5/40 + 4/40 = 17/40 of the job per hour
Therefore for the complete job to be done it will take 40/17 hours.

This may seem complex but let us give an example.
Let us say the job comprised of 40 units
Riva would do 8 units an hour for 5 hours
Megan would do 5 units an hour for 8 hours
Sarah would do 4 units and hour for 10 hours
Their combined effort per hour is 17 units
To do 40 units will therefore take 40/17 hours.